# Shahriar Noroozizadeh

### v1

@@ Appendix A Detailed background on the path-star task
A path-star graph $\mathcal{G}=({V},E)$ is a special tree graph from Bachmann and Nagarajan [12] that has a central node named ${v}_{\tt{root}}$ with multiple paths emanating from it. One of the leaf nodes is specified as ${v}_{\tt{goal}}$. The task is to find the unique path from ${v}_{\tt{root}}$ to ${v}_{\tt{goal}}$. To solve this task, one may either plan—by searching over the paths and backtracking—or execute a much simpler solution by following the unique path back from ${v}_{\tt{goal}}$, and then outputting the reverse of this path. Although this solution is algorithmically straightforward, and although a Transformer can even be shown to learn this solution under a simple multi-token modification to the objective, the next-token objective itself fails to learn this task even with sufficient amounts of data. Thus, the failure comes from optimization under the next-token objective. In this sense, the path-star task is known to be a minimal textbook adversarial example to next-token learning.
In-context path-star task. In the in-context version of the task, the model is given the prefix $\bm{p}=(\texttt{adj}(\mathcal{G}),{v}_{\tt{root}},{v}_{\tt{goal}})$ that provides a randomized adjacency list, and indicates the start and goal nodes, in the model’s context. The model must then produce the true path as response, $\bm{r}=({v}_{\tt{root}},\ldots,{v}_{\tt{goal}})$. To cast this as a learning task, we define a distribution $\mathcal{D}_{d,l}$ corresponding to graphs of the same topology, degree $d$ and path length $l$. To sample $\bm{p}\sim\mathcal{D}_{d,l}$, we uniformly sample node values from a vocabulary $\mathbb{V}$ to create the graph $G$; given this, we can randomize the adjacency list $\texttt{adj}(\mathcal{G})$. Figures˜3 and 2 contrast the two evaluation regimes we use throughout: in-weights (edge memorization & path finetuning) versus in-context (graph in the prompt).
@@ A.1 Failure of next-token learning in the in-context path-star task
Fig.˜5 demonstrates these failure modes empirically, showing that next-token prediction achieves only chance-level performance on path-star graphs of varying sizes, with the first token remaining particularly difficult to learn in isolation.
@@ B.2.2 Tiny model architectures
For quicker toy experiments on small-scale graphs (§C.3, §D.1.2), including the Tiny Path–Star, Tiny Grid, Tiny Cycle, and Tiny Irregular graphs, we used reduced-size architectures.
Models. We evaluated three model types: (1) a Transformer (TinyGPT), (2) a feed-forward neural network with the same configuration but without attention, and (3) a Mamba model of comparable scale to (1).
Configuration. For the Tiny Path–Star, Tiny Grid, and Tiny Cycle graphs, all models used a single layer ($N_{\tt layer}=1$), embedding dimension $m_{\tt width}=32$, and $m_{\tt head}=8$. The Mamba variant further included $d_{\text{state}}=8$, $d_{\text{conv}}=4$, and $\text{expand}=2$.
For the Tiny Irregular graphs, for all models, we used $N_{\tt layer}=3$, $m_{\tt width}=256$, and the Mamba parameters $d_{\text{state}}=8$, $d_{\text{conv}}=4$, $\text{expand}=4$. We use a larger number of layers since we found that otherwise the model does not achieve perfect edge-memorization with frozen embeddings.
Additional details. In associative-memory settings (e.g., left-column visualizations in Figure˜1), token embeddings were frozen. All models employed weight tying between input and output embeddings.
@@ B.5 Implementation and compute
Code is implemented in PyTorch with standard Transformer components. All runs fit on a single modern GPU (e.g., A100-40GB); per-figure training time depends on $(d,\ell)$ and batch size.
@@ C.1 Path-star task on Mamba SSM
This section demonstrates that the implicit reasoning on path-star graphs observed for Transformers in §2.1, generalizes to the Mamba SSM architecture [54] too. Fig.˜8 is the counterpart of the Transformer accuracy plots in Fig.˜4; Fig.˜9 the counterpart to Transformer heatmaps in Fig.˜16; Fig.˜10 the counterpart to the UMAP topology of Transformers in Fig.˜18. A notable difference here is that the Mamba model presents a strong geometry even when trained only on the edges (as seen in Fig.˜9(c)).
Figure 8: (left) Success of in-weights path-star task for Mamba. This figure is a counterpart to Fig.˜4. A next-token-trained Mamba achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜2.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜2.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first token) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.1.1). Success of this $\ell$-fold composition task is hard to explain within the associative memory (§2.2).
@@
Figure 9: Evidence of global geometry in path-star task for Mamba. This figure is a counterpart to Figs.˜6 and 16 for the Mamba SSM architecture. Recall that entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists. Interestingly, compared to the Transformer in Fig.˜16, the Mamba trained only on local supervision (right) exhibits a much stronger geometry.
Figure 10: UMAP projection of token embeddings of Mamba exhibits path-star topology. We corroborate the UMAP [98] observations from the Transformer (Fig.˜18) in Mamba SSM here. Each point is a node embedding; color indicates path identity. Different paths form separated clusters (although this clustering is weaker than for a Transformer); the central node ${v}_{\tt{root}}$ is excluded since it is shared across all paths. The graph has degree $10^{4}$ and path length 66 (including the root). Axes are UMAP components (arbitrary units). Note that we have re-used each color for multiple paths.
@@ C.2 Other large, harder path-finding graphs
In the tree-star graph, $\mathcal{T}_{d,\ell}$, there is a central node with degree $d$ and each child node except the leaf has a fixed degree of $2$, as visualized in Fig.˜11(a). The path-length from the root to any leaf is $\ell$. In this graph, two types of learning tasks can be considered, depending on how we split the test and training paths. For a no-overlap setting, we could sample all training paths from one subset of trees, and the test paths from the remaining trees; we call this the split at first token setting, since the test/train split is determined by the first token. A second setting—-with some test-train overlap—is one where we reserve some leaves as training goals, and the rest as test-goals. Here, on any test path, a prefix may have participated in a training path. We call this the split at leaf setting. Note that in both variants, all nodes will be sampled as part of the edge-memorization task.
In Fig.˜11(b), we find that on both variants the Transformer achieves non-trivial path-finding accuracy, generalizing our results beyond the path-star task. However, we note that the test-train split at the first token is much harder to succeed at, likely due to no overlaps.
@@ C.3 Tiny graphs
Besides the large path-star graph (of §2) and tree-star graphs (of C.2), we also report the embeddings learned on four tinier graphs for various architectures (details in B.2.2); the figures here are an extension of the Transformer embeddings in Fig.˜1. In these experiments, we train the models purely on local supervision (the edges, presented in both directions) to $100\%$ edge-memorization accuracy—for each vertex, we ensure that its $d$ neighbors appear in the top $d$ softmax probabilities. The graphs include (a) a tiny path-star graph with four paths of length $4$, (b) a $4\times 4$ grid graph, (c) a $15$-node cycle graph and (d) an irregular graph with two asymmetryic components. Each visualization includes the embeddings from: an associative memory model (implemented with a neural network with only one trainable matrix sandwiched between (un)embedding layers), a Node2Vec model, the eigenvectors of the graph Laplacian, a Transformer’s token embeddings, a neural network’s first layer, and a Mamba SSM’s token embeddings. For the Node2Vec model we use the top eigenvectors, and for the rest we use UMAP [98] to choose the top directions.
We consolidate the various observations from these figures (Figs.˜12, 13, 14 and 15) here:
In the tiny graphs of Figs.˜12, 13, 14 and 15, on various architectures (Node2Vec, graph spectrum, Transformer, neural network, Mamba SSM), we find that:
@@
    A geometry arises in all these architectures even without global supervision from a path-finding task (Section˜3.2).
@@
    The geometry arises in all three deep sequence models (Transformer, Mamba SSM, neural network) even though these models can learn the data associatively using the same learning setup, with just the (un)embedding matrices frozen (Section˜3.1).
@@
    The geometry of the Node2Vec model (which precludes associative memory), is much stronger than the deep sequence models, suggesting that the deep sequence models may be adulterated with associative memory (Section˜4).
@@
    We note that similar geometries arise even with only one direction presented; see §D.1.2. This is the scenario where both types of storage have the same bit and $\ell_{2}$ norm complexity; thus, there are no straightforward implicit pressures that encourage the geometry (Proposition˜1).

### v2

@@ Appendix A Detailed background on the path-star task
A path-star graph $\mathcal{G}=({V},E)$ is a special tree graph from Bachmann and Nagarajan [13] that has a central root named ${v}_{\tt{root}}$ with multiple disjoint, uniform-length paths emanating from it. One of the leaf nodes is specified as ${v}_{\tt{goal}}$. The task is to find the unique path from ${v}_{\tt{root}}$ to ${v}_{\tt{goal}}$. To solve this task, one may either plan—by searching over the paths and backtracking—or execute a much simpler solution by following the unique path back from ${v}_{\tt{goal}}$, and then outputting the reverse of this path. Although this solution is algorithmically straightforward, and although a Transformer can even be shown to learn this solution under a simple multi-token modification to the objective, the next-token objective itself fails to learn this task even with sufficient amounts of data. Thus, the failure comes from optimization under the next-token objective. In this sense, the path-star task is known to be a minimal textbook adversarial example to next-token learning.
In-context path-star task. In the in-context version of the task, the model is given the prefix $\bm{p}=(\texttt{adj}(\mathcal{G}),{v}_{\tt{root}},{v}_{\tt{goal}})$ that provides a randomized adjacency list, and indicates the start and goal nodes, in the model’s context. The model must then produce the true path as response, $\bm{r}=({v}_{\tt{root}},\ldots,{v}_{\tt{goal}})$. To cast this as a learning task, we define a distribution $\mathcal{D}_{d,l}$ corresponding to graphs of the same topology, degree $d$ and path length $l$. To sample $\bm{p}\sim\mathcal{D}_{d,l}$, we uniformly sample node values from a vocabulary $\mathbb{V}$ to create the graph $G$; given this, we can randomize the adjacency list $\texttt{adj}(\mathcal{G})$.
Figures˜3 and 2 contrast the two evaluation regimes we use throughout: in-weights (edge memorization & path finetuning) versus in-context (graph in the prompt).
@@ A.1 Failure of next-token learning in the in-context path-star task
Fig.˜5 (also reproduced in Fig.˜13) demonstrates these failure modes empirically, showing that next-token prediction achieves only chance-level performance on path-star graphs of varying sizes, with the first token remaining particularly difficult to learn in isolation.
(Success of implicit in-weights reasoning) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, trained on $75\%$ of the total $10^{4}$ paths, both the Transformer and Mamba are able to predict unseen paths when conditioned on held-out leaves with as much as 100% accuracy (see left plot of Fig.˜4 for Transformer (also reproduced in Fig.˜12), and left plot of Fig.˜14 for Mamba) Similar positive results for some harder graph topologies are in §C.2.
Shortly, we will isolate an even stronger instance of implicit reasoning from this task for analysis. To get there, let us scrutinize where the argument of B&N’24 may go differently in the in-weights setting for the model to succeed in Observation A.1.
A.1.1 Where does the path-star-failure argument go wrong in the in-weights setting?
(Model may experience future-token gradients) The in-weights path-star task is solved (Observation A.1) because Clever Hans cheats are not learned quickly enough, allowing future-token gradients to persist, teaching the model to find the right-to-left solution.
Figure 12: Success of Transformer in in-weights path-star task. (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜A.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜A.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.3). Analogous plots for Mamba are in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜13.
Figure 13: Failure of Transformer in in-context path-star task. (B&N’24) We report the failure of next-token Transformers in the in-context version of the path-star task, reproducing results from B&N’24. (left) Full path accuracy remains at chance level across different small graph sizes. (middle) Learning order of tokens with teacherless (multi-token-trained) objective shows a clear right-to-left learning cascade. (right) The hardest (first) token given the leaf fails to be learned in isolation, contrasting sharply with in-weights success shown in Fig.˜12.
(No reverse-learning cascade) The target tokens in the in-weights path-star task are learned in no particular order by the Transformer and Mamba models (see middle plots of Figs.˜4 and 14 in contrast with Fig.˜5; or see Fig.˜35 for side-by-side comparison).
The above observation weakens Section˜A.1.1, encouraging us to search for another one. A second possibility is that the second stage of the failure of B&N’24 does not trouble us in the in-weights task. Recall that in this stage, the first token is to be learned in isolation without future-token gradients. But this is an $\ell$-fold composition task that is theoretically well-understood to be computationally hard (as we elaborate later). Perhaps, that is not the case for our in-weights task:
Testing Section˜A.1.1 is easy: we train the model simply on the first token loss, instead of the loss over the full path sequence. We find that this task is trivial here, affirming Section˜A.1.1, and isolating a much stronger and cleaner instance of implicit in-weights reasoning.
@@ B.2.2 Tiny model architectures
For toy experiments on small-scale graphs (§C.3, §D.1.2), including the Tiny Path–Star, Tiny Grid, Tiny Cycle, and Tiny Irregular graphs, we used reduced-size architectures.
Models. We evaluated three model types: (1) a Transformer (TinyGPT), (2) a feed-forward neural network with the same configuration but without residual connection or attention and (TinyNN) (3) a Mamba model (TinySSM) of comparable scale to (1). Exact configurations are described below.
Configuration.
  • For all graphs, all models used a single layer ($N_{\tt layer}=1$), but the model specifications differed across graph types.
For TinyGPT and TinySSM in particular:
  • For Tiny Path–Star and Tiny Grid, TinyGPT used embedding dimension $m_{\tt width}=32$ with $m_{\tt head}=8$, and TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=2$, and $\text{expand}=1$.
  • For Tiny Cycle, TinyGPT used embedding dimension $16$ with $8$ heads, while TinySSM matched the same specifications as above ($8$-dim, $d_{\text{state}}=2$, $d_{\text{conv}}=2$, $\text{expand}=1$).
  • For Tiny Irregular, TinyGPT again used embedding dimension $16$ with $8$ heads, and TinySSM used embedding dimension $8$ with $d_{\text{state}}=2$, $d_{\text{conv}}=4$, and $\text{expand}=1$.
For TinyNN, we used a large embedding dimension $512$ with $1$ head for the sake of empirical proofs of Section˜3. We elaborate why. First, TinyNN architectures are a clean architecture where one can be certain that freezing the (un)embedding layers results in a purely associative storage; thus they are appropriate for our empirical proofs. Next, in our proofs, we wanted to make two demonstrations. First, that the model memorize geometrically even without bottleneck pressures (as is the case with a large embedding dimension); second, that in sufficiently wide-neck settings, the associative storage is realized within just $2$ learning steps (which we find is true here).
We clarify that even when reducing the embedding dimension of TinyNN to match that of TinyGPT (Tiny Path–Star and Tiny Grid: $m_{\tt width}=32$; Tiny Cycle and Tiny Irregular: $m_{\tt width}=16$), we found that the geometric structure still persisted and that the associative memory variant continued to fully memorize the edge bigrams, albeit with longer optimization steps.
Associative memory variants of these architectures. In associative-memory settings (e.g., left-column visualizations in Figure˜1), token embeddings were frozen in all these architectures. All models employed weight tying between input and output embeddings. Note that since we employ only one trainable layer in these settings (the embedding/unembedding layers frozen), we can be certain, at least in the case of TinyNN that storage is associative.
@@ B.5 Implementation and compute
Code is implemented in PyTorch with standard Transformer components. All runs fit on a single modern GPU (e.g., A100-40GB or A100-80GB for the largest batch-size); per-figure training time depends on $(d,\ell)$ and batch size.
@@ C.1 Path-star task on Mamba SSM
This section demonstrates that the implicit reasoning on path-star graphs observed for Transformers in §2.1, generalizes to the Mamba SSM architecture [60] too. Fig.˜14 is the counterpart of the Transformer accuracy plots in Fig.˜4; Fig.˜15 is the counterpart to Transformer heatmaps in Fig.˜25; Fig.˜16 is the counterpart to the UMAP topology of Transformers in Fig.˜27. A notable difference here is that the Mamba model presents a strong geometry even when trained only on the edges (as seen in Fig.˜15(c)).
Figure 14: (left) Success of in-weights path-star task for Mamba. This figure is a counterpart to Fig.˜4. A next-token-trained Mamba achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜A.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜A.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first token) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory (§2.3).
@@
Figure 15: Evidence of global geometry in path-star task for Mamba. This figure is a counterpart to Figs.˜6 and 25 for the Mamba SSM architecture. Recall that entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists. Interestingly, compared to the Transformer in Fig.˜25, the Mamba trained only on local supervision (right) exhibits a much stronger geometry.
Figure 16: UMAP projection of token embeddings of Mamba exhibits path-star topology. We corroborate the UMAP [109] observations from the Transformer (Fig.˜27) in Mamba SSM here. Each point is a node embedding; color indicates path identity. Different paths form separated clusters (although this clustering is weaker than for a Transformer); the central node ${v}_{\tt{root}}$ is excluded since it is shared across all paths. The graph has degree $10^{4}$ and path length 66 (including the root). Axes are UMAP components (arbitrary units). Note that we have re-used each color for multiple paths.
@@ C.2 Other large, harder path-finding graphs
In the tree-star graph, $\mathcal{T}_{d,\ell}$, there is a central node with degree $d$ and each child node except the leaf has a fixed degree of $2$, as visualized in Fig.˜17(a). The path-length from the root to any leaf is $\ell$. In this graph, two types of learning tasks can be considered, depending on how we split the test and training paths. For a no-overlap setting, we could sample all training paths from one subset of trees, and the test paths from the remaining trees; we call this the split at first token setting, since the test/train split is determined by the first token. A second setting—with some test-train overlap—is one where we reserve some leaves as training goals, and the rest as test-goals. Here, on any test path, a prefix may have participated in a training path. We call this the split at leaf setting. Note that in both variants, all nodes will be sampled as part of the edge-memorization task.
In Fig.˜17(b), we find that on both variants the Transformer achieves non-trivial path-finding accuracy, generalizing our results beyond the path-star task. However, we note that the test-train split at the first token is much harder to succeed at, likely due to no overlaps.
@@ C.3 Tiny graphs
Besides the large path-star graph (of §2) and tree-star graphs (of C.2), we also report the embeddings learned on four tinier graphs for various architectures (details in B.2.2); the figures here are an extension of the Transformer embeddings in Fig.˜1. In these experiments, we train the models purely on local supervision (the edges, presented in both directions) to $100\%$ edge-memorization accuracy: for each vertex, we ensure that its $d$ neighbors appear in the top $d$ softmax probabilities. The graphs include (a) a tiny path-star graph with four paths of length $4$, (b) a $4\times 4$ grid graph, (c) a $15$-node cycle graph and (d) an irregular graph with two asymmetric components. Each visualization includes the embeddings from: an associative memory model (implemented with a neural network with only one trainable matrix sandwiched between (un)embedding layers), a Node2Vec model, the eigenvectors of the graph Laplacian, a Transformer’s token embeddings, a neural network’s first layer, and a Mamba SSM’s token embeddings. For the Node2Vec model we use the top eigenvectors, and for the rest we use UMAP [109] to choose the top directions.
We note that these visualizations by themselves do not guarantee the presence of global information (a model that embeds each node as its adjacency row can still show similar geometries). As a rigorous test, we provide node-node cosine-similarity heatmaps to examine whether it exhibits multi-hop information. These are in Figs.˜24 and 23.
We consolidate the various observations from these figures (Figs.˜18, 19, 20, 21, 24, 23 and 22) and our memorization vs. time plots in Fig.˜7 here:
In the tiny graphs of Figs.˜18, 19, 20 and 21, on various architectures (Node2Vec, graph spectrum, Transformer, neural network, Mamba SSM), we find that:
@@
    A geometry arises in all these architectures even without global supervision from a path-finding task (Section˜3.1).
@@
    The geometry arises in all three deep sequence models (Transformer, Mamba SSM, neural network) even though these models can learn the data associatively using the same learning setup, with just the (un)embedding matrices frozen (Section˜3.2).
@@
    The geometry of the Node2Vec model (which precludes associative memory), is much stronger than the deep sequence models, suggesting that the deep sequence models may be adulterated with associative memory (˜1).
@@
    Associative storage can be quickly found by gradient descent (Fig.˜7), as claimed in Section˜3.2, whereas geometric memorization takes longer to discover (Fig.˜22)
  6. 6.
    We note that similar geometries arise even with only one direction presented; see §D.1.2. This is the scenario where both types of storage have the same bit and $\ell_{2}$ norm complexity; thus, there are no straightforward implicit pressures that encourage the geometry (Section˜3.2).
@@
(a) Tiny Path-Star (LR=$0.001$)
(b) Tiny Path-Star (LR=$0.01$)
(c) Tiny Path-Star (LR=$0.1$)
(d) Tiny Grid (LR=$0.001$)
(e) Tiny Grid (LR=$0.01$)
(f) Tiny Grid (LR=$0.1$)
(g) Tiny Cycle (LR=$0.001$)
(h) Tiny Cycle (LR=$0.01$)
(i) Tiny Cycle (LR=$0.1$)
(j) Tiny Irregular (LR=$0.001$)
(k) Tiny Irregular (LR=$0.01$)
(l) Tiny Irregular (LR=$0.1$)
Figure 22: For various learning rates, geometric memorization takes much longer for gradient descent to discover: This is an extended version of Fig.˜8 for two more learning rates. Recall that for the same TinyNN architecture as in Fig.˜7 but without frozen embeddings, we roughly quantify geometric memorization over time via cosine distance across all pairs of vertices, alongside 2D visualizations. Each row is a different graph, and each column is a different learning rate. For both the learning rate $0.001$ and $0.01$, the geometries take about $200$ steps to appear; the larger learning rate of $0.1$ is too aggressive to create the geometry. Recall that, for associative memory, two steps with learning rate of $0.1$ sufficed (Fig.˜7). This is evidence for Section˜3.2: ease of discovery does not determine what sort of memory is preferred.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 23: Node-node cosine similarity vs. adjacency matrix for the Transformer: For each graph, we plot the cosine similarity between all node embeddings and compare it against the adjacency matrix. Observe that the cosine similarities exhibit a richer structure than the adjacency matrix, reflecting some notion of multi-hop distance e.g., in the cycle graph, there is a gradual decrease in similarity as we walk towards the diametrically opposite node in any given row.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 24: Node-node cosine similarity vs. adjacency matrix for the Node2Vec model: Like in the Transformer (Fig.˜23), we again observe that cosine similarities exhibit a rich, multi-hop structure absent in the adjacency matrix.